OpenStudy (anonymous):

rewrite as a base power, if possible. x^3 / x

OpenStudy (anonymous):

The x has a power of 1

OpenStudy (anonymous):

So (x^3)/(x^1) = x^2.

OpenStudy (anonymous):

(x *x *x) -------- x The two x's on the left most side cancel. Leaving you (x*x)/1 = x^2

OpenStudy (anonymous):

thank you